OpenStudy (anonymous):

Solve. \[7^{x-4} = 7 ^{4x+7}\]

myininaya (myininaya):

just set the exponents equal and solve for x x-4=4x+7

OpenStudy (anonymous):

i get -11/3 and thats not one of the choices.

myininaya (myininaya):

Subtract x on both sides -4=3x+7 Subtract 7 on both sides -11=3x and divide by 3 on both sides -11/3=x I agree with you answer Is it perhaps written in decimal form or as a mix fraction?

myininaya (myininaya):

\[x=-\frac{3(3)+2}{3}=-(\frac{3(3)}{3}+\frac{2}{3})=-(3+\frac{2}{3})=-3\frac {2}{3}\]

OpenStudy (anonymous):

the answers are A: x = -4, B: x=2/3, C:x=-12/5, D: x=4

myininaya (myininaya):

The question is correct, right?

OpenStudy (anonymous):

it is \[-\frac{11}{3}\] for sure

myininaya (myininaya):

yep unless the problem is wrong